For this problem write the equation that expresses the relationship among the variables, and then use the given data to solve for the variation constant


y is directly proportional to x, and y=38 when x=114

Answer by Theo(2978)

y = k*x

if y = 38 and x = 114, this becomes:

38 = k*114

k = 38/114

k = (1/3)

This is saying that y = (1/3) * x

38 = (1/3) * 114

38 = 38 confirming the value for k is good.

they are directly proportional which means that an increase in one yields an increase in the other by the stated proportion.

suppose y = 1

formula states that 1 = (1/3)*x which makes x = 3.

An increase of 1 in y will result in an increase of 3 in x.

2 = (1/3)*x makes x = 6

y increased by 1
x increased by 3